a cylindrical water tower has a radius of 10ft. and a height of twelve feet. what is the weigth of the water tower when it is full if one cubic foot of water weighs8.1 pounds
Answer by checkley75(3666)

Volume of this tower is: area of base=2piR
base=2*3.14*10
base=6.28*10
base=62.8 ft^2
volume=height*base
volume=12*62.8
volume=753.6 ft^3 is the volume;
753.6*8.1=6104.16 pounds is the weight of the water in the tower.